OpenStudy (anonymous):

who can help explain how to solve this problem for me... 5(31/16)+3y=-11

OpenStudy (anonymous):

ok first think you are going to want to do is clear the fraction because, well lets face it nobody likes them.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

to do this you are going to multiply everything by 16 so that you end up with \[31+48y=-176\] no we have a much easier number to work with

OpenStudy (anonymous):

the next step is to get your varibles on one side and constants on the other

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

retype in the equation please

OpenStudy (anonymous):

48y=-176-31?

OpenStudy (anonymous):

was the first number 5 and 31/16 ?

OpenStudy (anonymous):

5 times 31/15

OpenStudy (anonymous):

ok thats going to change up the numbers a bit

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

if you can please explain to me i have finals next week and a exam friday i want to really understand this!

OpenStudy (anonymous):

but then you after you have transposed your constants and vvaribles you divide both sides by the coeficent of the varible to get your answer

OpenStudy (anonymous):

im really confused now?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

x=-331/48 try to simplify

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

5(31/16)+3y=-11 => 155+48y=-176 =>48y=-176-155 => 48y=-331 => y=-331/48

OpenStudy (anonymous):

ok let me see if i get it right

OpenStudy (anonymous):

it didnt come out right i think i did it wrong

OpenStudy (anonymous):

it said to solve using the elimination method and it gave me 2 diffrent problems

OpenStudy (anonymous):

5x+3y=-11 7x-2y=19 solve using elimination method

OpenStudy (anonymous):

can u help me with this?

OpenStudy (anonymous):

multiply the top function by 2 and the bottom function by 3 and tell me what you get

OpenStudy (anonymous):

10x+6y=-22 21x-6y=38

OpenStudy (anonymous):

then i would get 31x=16

OpenStudy (anonymous):

ok now you are going to combine the top and bottom function, and yes

OpenStudy (anonymous):

so then i would get x= 31/16

OpenStudy (anonymous):

after that i dont know what to do

OpenStudy (anonymous):

it would be 16/31

OpenStudy (anonymous):

sub that in for your x value inorder to get your y value

OpenStudy (anonymous):

so i would do 5(31/16)+3y=-11

OpenStudy (anonymous):

5 times 31/16

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

and u said y=-331/48

OpenStudy (anonymous):

and we just did that problem

OpenStudy (anonymous):

and when i put it in my homework it said that was wrong

OpenStudy (anonymous):

hmmm...

OpenStudy (anonymous):

i put (31/16,-331/48)

OpenStudy (anonymous):

as the two points

OpenStudy (anonymous):

did you normally enter them in as improper fractions or mixed numbers?

OpenStudy (anonymous):

it said the correct answer was (35/31,172/31)

OpenStudy (anonymous):

and i dont understand how they get that answer?!

OpenStudy (anonymous):

I don't either i am sorry i couldn't help

OpenStudy (anonymous):

its ok thanks anyway :)